# An Operator Splitting Method for Large-Scale CVaR-Constrained Quadratic Programs

### v3

@@ 1 Introduction
Many applications in science and engineering require optimizing expected performance while controlling the risk of extreme outcomes. The conditional value-at-risk (CVaR) is one of the most widely used risk measures in practice. It measures the conditional expectation of losses exceeding a given threshold, effectively capturing the severity of tail events. Unlike variance-based measures, CVaR penalizes only downside risk, aligning with how practitioners view risk in real applications. It provides a clear interpretation as the expected value in the tail of the loss distribution and satisfies the mathematical properties of a coherent risk measure: monotonicity, sub-additivity, homogeneity, and translation invariance. Most importantly for optimization, CVaR is convex, which means the resulting optimization problems are tractable and can be solved reliably and efficiently.
@@
We present two main contributions. First, we develop a specialized $O(m\log m)$ algorithm for projecting onto CVaR constraints, where $m$ is the number of scenarios. Building on this algorithm, our second contribution is an operator splitting method for solving large-scale CVaR-constrained quadratic programs. The method alternates between solving a linear system and performing parallel projections: one onto the CVaR constraints using our specialized algorithm and another onto box constraints with a simple closed-form solution. Numerical examples from several application domains demonstrate that our method outperforms general-purpose solvers by several orders of magnitude on problems with up to millions of scenarios.
@@ 1.1 Conditional value-at-risk (CVaR)
For a finite set of samples $z_{1},\ldots,z_{m}\in{\mbox{\bf R}}$ representing the distribution of losses, Rockafellar and Uryasev [rockafellar2000optimization] showed that the CVaR at level $\beta$, denoted $\phi_{\beta}:{\mbox{\bf R}}^{m}\to{\mbox{\bf R}}$, can be computed as
@@ 1.2 CVaR-constrained quadratic programs
where $x\in{\mbox{\bf R}}^{n}$ is the decision variable. The objective function is defined by a positive semidefinite matrix $P\in{\mbox{\bf S}}^{n}_{+}$ and a vector $q\in{\mbox{\bf R}}^{n}$. The CVaR constraint is defined by quantile level $\beta\in(0,1)$, threshold $\kappa\in{\mbox{\bf R}}$, and matrix $A\in{\mbox{\bf R}}^{m\times n}$. Additional linear inequality constraints are defined by a matrix $B\in{\mbox{\bf R}}^{p\times n}$, and vectors $l\in\left({{\mbox{\bf R}}\cup\{-\infty\}}\right)^{p}$, $u\in\left({{\mbox{\bf R}}\cup\{\infty\}}\right)^{p}$. We focus on the case where $p\ll m$, i.e., the number of additional constraints is much smaller than the number of scenarios.
We assume the CVQP problem (1) is feasible and that $P$, $A$, and $B$ have zero common nullspace, which is sufficient to ensure that the problem is bounded.
@@
With $z=(x,t)$, this matches our standard CVQP form (1).
@@
The CVQP is a convex optimization problem, and by using the linear inequality representation of the CVaR constraint, it can be represented as a quadratic program. As such, the problem can be solved using standard solvers that support quadratic programs. However, when the number of samples (sometimes called scenarios) $m$ is very large, solving the problem as a generic quadratic program can become prohibitively slow.
@@
CVaR-constrained quadratic programs appear in many domains. Portfolio optimization uses CVaR constraints to bound worst-case losses [krokhmal2002portfolio]. CVaR constraints are also used to manage tail risk in supply chains [sawik2011selection], network flows [sorokin2013computational], and energy systems [doege2006risk]. In control and statistical learning, CVaR provides convex approximations for chance constraints [van2015distributionally, laguel2021superquantiles]. For a comprehensive review, see [filippi2020conditional].
@@ 1.3 Related work
CVaR constraints were introduced by Rockafellar and Uryasev [rockafellar2000optimization] and are now standard in risk-constrained optimization [krokhmal2002portfolio]. Our method builds on operator splitting [boyd2011distributed], which decomposes the problem into simpler subproblems: a linear system solve and projections.
The CVaR projection algorithm adds to a family of finite-termination projection methods, including the classical simplex projection [gafni1984two] and isotonic regression [barlow1972isotonic]. Concurrently with the development of our CVaR projection algorithm, Roth and Cui released a preprint [roth2023n], later published as [roth2025n], containing two finite termination algorithms for CVaR projection attaining the same complexity. After corresponding with the authors, we were unable to find any equivalence between their algorithms and ours.
Roth and Cui [roth2024fast] also proposed a second-order computational framework for CVaR-constrained optimization. Their approach differs from our first-order ADMM method, instead using a semismooth Newton-based augmented Lagrangian method.
@@ 2.1 ADMM
We will solve the CVQP by reformulating the problem and then using the alternating direction method of multipliers (ADMM). By introducing auxiliary variables $z\in{\mbox{\bf R}}^{m}$ and $\tilde{z}\in{\mbox{\bf R}}^{p}$, we can rewrite problem (1) as
$\begin{array}[]{ll}\mbox{minimize}&(1/2)x^{T}Px+q^{T}x+I_{\mathcal{C}}(z)+I_{[l,u]}(\tilde{z})\\ \mbox{subject to}&Ax=z,\quad Bx=\tilde{z},\end{array}$
where $\mathcal{C}=\{z\mid\phi_{\beta}(z)\leq\kappa\}$. Here, $I_{\mathcal{C}}$ and $I_{[l,u]}$ are the indicator functions for the sets $\mathcal{C}$ and $\{\tilde{z}\mid l\leq\tilde{z}\leq u\}$, respectively, given by
@@
We use the scaled dual variables $u=(1/\rho)y$ and $\tilde{u}=(1/\rho)\tilde{y}$, where $\rho>0$ is a hyperparameter and $y\in{\mbox{\bf R}}^{m}$ and $\tilde{y}\in{\mbox{\bf R}}^{p}$ are the vectors of dual variables associated with the equality constraints $Ax=z$ and $Bx=\tilde{z}$, respectively. We use ADMM with over-relaxation (see [eckstein1992douglas, eckstein1994parallel] for analysis), with over-relaxation parameter $\alpha\in(0,2)$.
@@
Our assumption on the common nullspace of $P$, $A$, and $B$ implies that $M$ is positive definite. The ADMM updates are
$x^{k+1}$ ::=:= $\mathop{\rm argmin}_{x}\left(\frac{1}{2}x^{T}Mx+(p^{k})^{T}x\right)$ (2)
$z^{k+1/2}$ ::=:= $\alpha Ax^{k+1}+(1-\alpha)z^{k}$ (3)
$\tilde{z}^{k+1/2}$ ::=:= $\alpha Bx^{k+1}+(1-\alpha)\tilde{z}^{k}$ (4)
$z^{k+1}$ ::=:= $\Pi_{\mathcal{C}}\left(z^{k+1/2}+u^{k}\right)$ (5)
$\tilde{z}^{k+1}$ ::=:= $\Pi_{[l,u]}\left(\tilde{z}^{k+1/2}+\tilde{u}^{k}\right)$ (6)
$u^{k+1}$ ::=:= $u^{k}+z^{k+1/2}-z^{k+1}$ (7)
$\tilde{u}^{k+1}$ ::=:= $\tilde{u}^{k}+\tilde{z}^{k+1/2}-\tilde{z}^{k+1}.$ (8)
The operator $\Pi_{\mathcal{C}}$ is the (Euclidean) projection onto $\mathcal{C}$ and $\Pi_{[l,u]}$ is the projection onto the set $[l,u]\subset{\mbox{\bf R}}^{p}$.
It is well known that this ADMM algorithm converges [eckstein1992douglas, GABAY1983299].
@@ 2.2 Evaluating the updates
That leaves only the update (5), projection onto the CVaR constraint set. In the next section we present our core contribution, an efficient algorithm for evaluating this projection operator. The algorithm we present has a complexity of $O(m\log m)$, where $m$ is the number of scenarios in the CVaR constraint.
@@ 3 CVaR projection preliminaries
We now present preliminaries which motivate our efficient algorithm for evaluating the projection operator $\Pi_{\mathcal{C}}$, i.e., for solving the problem
@@
where $z_{[1]}\geq z_{[2]}\geq\cdots\geq z_{[m]}$ are the components of $z$ in nonincreasing order. This is a convex function that sums the $k$ largest elements of a vector. We rewrite problem (9) as
@@
where $k=(1-\beta)m$ and $d=\kappa k$. We assume $(1-\beta)m\in{\mbox{\bf N}}$ (and round up to the nearest integer if not).
@@ 3.1 Sorted projection
Here $u$ is the permuted change of variable $u=Pz$.
We can sort $v$ in descending order via $P$, project the sorted vector $v^{\prime}$, and unsort the result via $P^{T}$ to obtain the projection of $v$ onto the set $\{z\mid f_{k}(z)\leq d\}$.
@@ 3.2 Optimality conditions
We characterize the optimality conditions of the projection problem; these conditions will be used in §4.3 to prove the correctness of the algorithm. We can rewrite the problem (10) in an equivalent form as
$\begin{array}[]{ll}\mbox{minimize}&\frac{1}{2}\|v-z\|_{2}^{2}\\ \mbox{subject to}&a_{i}^{T}z\leq d,\quad i=1,\ldots,\binom{m}{k},\end{array}$
the subscript $i$ is an enumeration of
$\mathcal{A}=\left\{a\in\{0,1\}^{m}\mid\mathbf{1}^{T}a=k\right\},$
the set of all $\binom{m}{k}$ $k$-hot indicator vectors.
@@
The optimality conditions motivate an algorithm which solves the KKT system. We will define $z=v-\Delta$, where $\Delta$ is the vector we remove from $v$ to get the projection. The KKT conditions above imply that $\Delta$ is a nonnegative combination of the vectors $a_{i}$ for which $a_{i}^{T}z=d$, i.e., the indicators of the largest $k$ entries of $z$. (Note that there are potentially several such vectors in the presence of ties: consider the largest 2 entries in the vector $(2,1,1,0)$.) Thus, if we can construct a vector $\Delta$ which is a nonnegative combination of largest-$k$ indicator vectors, such that $f_{k}(v-\Delta)=d$, then we have found the projection $z=v-\Delta$.
@@ 4.1 Decrease step
is the set of $k$-hot indicator vectors who have the first $n_{u}$ entries equal to $1$ and all $1^{\prime}s$ occurring in the first $n_{u}+n_{t}$ entries.
@@ 4.2 Complexity
At each step, $n_{u}$ either decreases by $1$, or remains the same. At each step, $n_{t}$ increase by $1$. If the algorithm hasn’t terminated after $m$ steps, then there are $m$ tied entries, thus the subsequent decrease step will reduce the sum to the desired value, and the algorithm will terminate. Each step of the algorithm has constant complexity, and thus the algorithm (with sorted input) has complexity $O(m)$. Therefore, including the sorting and unsorting, the total complexity of the algorithm is $O(m\log m)$.
@@ 5 Numerical experiments
In this section, we compare our methods against two state-of-the-art solvers: Mosek (commercial) and Clarabel (open-source). We benchmark our specialized CVaR projection algorithm on random problem instances, then evaluate our ADMM-based method on large-scale CVQPs from portfolio optimization and quantile regression. Code is available at https://github.com/cvxgrp/cvqp.
@@ 5.1 Experimental setup
We ran all experiments on a Google Cloud n1-highmem-32 instance (32 vCPUs, 208GB memory, Intel Haswell platform) with a 2-hour time limit for benchmarks.
For our CVQP solver, we set the initial penalty parameter $\rho=10^{-2}$ and over-relaxation parameter $\alpha=1.7$. We update $\rho$ adaptively: when the ratio of primal to dual residual exceeds $\mu=10$, we scale $\rho$ by $\tau=2$ in the appropriate direction. The algorithm stops when both primal and dual residuals fall below $10^{-4}$ absolute and $10^{-3}$ relative tolerance.
@@ 5.2 CVaR projection
We generated random CVaR projection problems with vectors $v\in{\mbox{\bf R}}^{m}$ having entries uniformly distributed on $[0,1]$. For each $v$, we set the constraint threshold to $d=\tau f_{k}(v)$ in problem (11), where $\tau\in(0,1)$ controls problem difficulty (larger $\tau$ gives easier problems). The task is to project $v$ onto the convex set $\{z\mid f_{k}(z)\leq d\}$.

### v4

@@ 1 Introduction
Many applications in science and engineering require controlling the risk of extreme outcomes. The conditional value-at-risk (CVaR) is the expected value of losses exceeding a given quantile. Unlike variance, it penalizes only downside risk; unlike value-at-risk, it is coherent and convex. Convexity means that optimization problems involving CVaR can be solved exactly and efficiently.
@@
We present two main contributions. First, we develop an $O(m\log m)$ algorithm for projecting onto CVaR constraints, where $m$ is the number of scenarios. Building on this algorithm, our second contribution is an operator splitting method for solving large-scale CVaR-constrained quadratic programs. The method alternates between solving a linear system and performing parallel projections: one onto the CVaR constraints using our specialized algorithm and another onto box constraints with a simple closed-form solution. Numerical examples from several application domains demonstrate that our method outperforms general-purpose solvers by several orders of magnitude on problems with up to millions of scenarios.
@@ 1.1 Conditional value-at-risk (CVaR)
For a finite set of samples $z_{1},\ldots,z_{m}\in{\mbox{\bf R}}$ representing the distribution of losses, Rockafellar and Uryasev [13] showed that the CVaR at level $\beta$, denoted $\phi_{\beta}:{\mbox{\bf R}}^{m}\to{\mbox{\bf R}}$, can be computed as
@@ 1.2 CVaR-constrained quadratic programs
where $x\in{\mbox{\bf R}}^{n}$ is the decision variable. The objective function is defined by a positive semidefinite matrix $P\in{\mbox{\bf S}}^{n}_{+}$ and a vector $q\in{\mbox{\bf R}}^{n}$. The CVaR constraint is defined by quantile level $\beta\in(0,1)$, threshold $\kappa\in{\mbox{\bf R}}$, and matrix $A\in{\mbox{\bf R}}^{m\times n}$. Additional linear inequality constraints are defined by a matrix $B\in{\mbox{\bf R}}^{p\times n}$, and vectors $l\in\left({{\mbox{\bf R}}\cup\{-\infty\}}\right)^{p}$, $u\in\left({{\mbox{\bf R}}\cup\{\infty\}}\right)^{p}$. Since $P$ is positive semidefinite and the constraints are convex, the CVQP is a convex optimization problem. We focus on the case where $p\ll m$, i.e., the number of additional constraints is much smaller than the number of scenarios.
We assume the CVQP problem (1) is feasible and that $P$, AA, and $B$ have zero common nullspace, which ensures that the problem is bounded. This condition also implies that the matrix $M=P+\rho(A^{T}A+B^{T}B)$ is positive definite for any $\rho>0$, which is needed for the ADMM linear system solve described in §2. In practice the condition holds when $P$ is positive definite or when $A$ has full column rank.
@@
Defining the decision variable as the pair $(x,t)$, this is an instance of the standard CVQP form (1).
@@
By using the linear inequality representation of the CVaR constraint, the CVQP can be reformulated as a standard quadratic program and solved with any QP solver. However, when the number of scenarios $m$ is large, this approach becomes prohibitively slow.
@@
CVaR-constrained quadratic programs arise in many domains. In portfolio optimization, $x$ represents portfolio weights and $Ax$ gives portfolio losses across $m$ return scenarios; the CVaR constraint limits the expected loss in the worst $(1-\beta)$ fraction of scenarios, while the quadratic objective captures risk-adjusted returns [10]. In supply chain planning, $x$ includes order quantities and routing decisions, with scenarios modeling demand uncertainty and supplier disruptions; CVaR constraints bound extreme total cost realizations [18, 9]. Similar formulations appear in network flow problems with stochastic capacities [19], facility location and disaster response logistics [12], energy systems with uncertain generation and prices [3], and radiation treatment planning, where CVaR constraints replace intractable dose-volume constraints [14]. In control and statistical learning, CVaR provides tractable convex approximations to chance constraints [20, 11]. For a comprehensive review, see [6].
@@ 1.3 Related work
CVaR constraints were introduced by Rockafellar and Uryasev [13] and are now standard in risk-constrained optimization [10]. Our method builds on operator splitting [2], which decomposes the problem into simpler subproblems: a linear system solve and projections.
The CVaR projection algorithm adds to a family of finite-termination projection methods, including the classical simplex projection [8] and isotonic regression [1]. Concurrently with the development of our CVaR projection algorithm, Roth and Cui released a preprint [15], later published as [17], containing two finite termination algorithms for CVaR projection attaining the same complexity. After corresponding with the authors, we were unable to find any equivalence between their algorithms and ours.
Roth and Cui [16] also proposed a second-order computational framework for CVaR-constrained optimization, using a semismooth Newton-based augmented Lagrangian method. As a second-order method, it can achieve superlinear convergence near a solution, while ADMM is a first-order method with an $O(1/k)$ convergence rate. On the other hand, each ADMM iteration is cheap (a single backsolve and parallel projections after an initial factorization), and the method is simple to implement. In our experience, ADMM reaches moderate accuracy in few iterations, which is sufficient for most applications.
@@ 2.1 ADMM
We solve the CVQP by reformulating the problem and applying the alternating direction method of multipliers (ADMM). By introducing auxiliary variables $z\in{\mbox{\bf R}}^{m}$ and $\tilde{z}\in{\mbox{\bf R}}^{p}$, we can rewrite problem (1) as
[…]
where $\mathcal{C}=\{z\mid\phi_{\beta}(z)\leq\kappa\}$. The auxiliary variables $z$ and $\tilde{z}$ decouple the CVaR constraint from the box constraints, so that each can be handled separately in the ADMM updates below. Here, $I_{\mathcal{C}}$ and $I_{[l,u]}$ are the indicator functions for the sets $\mathcal{C}$ and $\{\tilde{z}\mid l\leq\tilde{z}\leq u\}$, respectively, given by
@@
We use the scaled dual variables $u=(1/\rho)y$ and $\tilde{u}=(1/\rho)\tilde{y}$, where $\rho>0$ is a hyperparameter and $y\in{\mbox{\bf R}}^{m}$ and $\tilde{y}\in{\mbox{\bf R}}^{p}$ are the vectors of dual variables associated with the equality constraints $Ax=z$ and $Bx=\tilde{z}$, respectively. We use ADMM with over-relaxation (see [4, 5] for analysis), with over-relaxation parameter $\alpha\in(0,2)$.
@@
Our assumption on the common nullspace of $P$, AA, and $B$ implies that $M$ is positive definite. The ADMM updates are
$x^{k+1}$ $:=$ $\mathop{\rm argmin}_{x}\left(\frac{1}{2}x^{T}Mx+(p^{k})^{T}x\right)$ (2)
$z^{k+1/2}$ $:=$ $\alpha Ax^{k+1}+(1-\alpha)z^{k}$ (3)
$\tilde{z}^{k+1/2}$ $:=$ $\alpha Bx^{k+1}+(1-\alpha)\tilde{z}^{k}$ (4)
$z^{k+1}$ $:=$ $\Pi_{\mathcal{C}}\left(z^{k+1/2}+u^{k}\right)$ (5)
$\tilde{z}^{k+1}$ $:=$ $\Pi_{[l,u]}\left(\tilde{z}^{k+1/2}+\tilde{u}^{k}\right)$ (6)
$u^{k+1}$ $:=$ $u^{k}+z^{k+1/2}-z^{k+1}$ (7)
$\tilde{u}^{k+1}$ $:=$ $\tilde{u}^{k}+\tilde{z}^{k+1/2}-\tilde{z}^{k+1}.$ (8)
The operator $\Pi_{\mathcal{C}}$ is the (Euclidean) projection onto $\mathcal{C}$ and $\Pi_{[l,u]}$ is the projection onto the set $l,u]\subset{\mbox{\bf R}}^{p}$. The over-relaxation, dual updates, and box projection are all trivial operations; the linear system solve requires one cached factorization of $M$. The main computational bottleneck is the CVaR projection (5), which we address in §34.
It is well known that this ADMM algorithm converges [4, 7], with a worst-case $O(1/k)$ convergence rate, though linear convergence is commonly observed in practice for quadratic programs [2].
@@ 2.2 Evaluating the updates
That leaves only the CVaR projection (5), for which we develop an $O(m\log m)$ algorithm in the next two sections.
2.3 Summary
For convenience we summarize the complete CVQP method in Algorithm 1. The CVaR projection is the main computational bottleneck and uses the $O(m\log m)$ algorithm described in §4.
Algorithm 1 CVQP solver
1:Factorize $M=P+\rho(A^{T}A+B^{T}B)$.
2:repeat
3:  Solve linear system (2).
4:  Over-relax (3)–(4).
5:  Project onto CVaR constraint (5); see §4.
6:  Clip to box constraints (6).
7:  Update dual variables (7)–(8).
8:  Update $\rho$ and re-factorize $M$ if needed.
9:until primal and dual residuals below tolerance.
@@ 3 CVaR projection preliminaries
We now present preliminaries that motivate our efficient algorithm for evaluating the projection operator $\Pi_{\mathcal{C}}$, i.e., for solving the problem
@@
where $z_{[1]}\geq z_{[2]}\geq\cdots\geq z_{[m]}$ are the components of $z$ in nonincreasing order. This is a convex function that sums the $k$ largest elements of a vector (the pointwise maximum of $\binom{m}{k}$ linear functions). Setting $k=(1-\beta)m$ and evaluating the infimum over $\alpha$ in the Rockafellar-Uryasev formula (the minimizer is $\alpha=z_{[k]}$), we obtain $\phi_{\beta}(z)=(1/k)f_{k}(z)$. The constraint $\phi_{\beta}(z)\leq\kappa$ is therefore equivalent to $f_{k}(z)\leq\kappa k$. We rewrite problem (9) as
@@
where $d=\kappa k$. We assume $(1-\beta)m\in{\mbox{\bf N}}$ (and round up to the nearest integer if not).
@@ 3.1 Sorted projection
The last equality is the change of variable $u=Pz$. In words, we can sort $v$ in descending order, project the sorted vector $v^{\prime}$, and unsort the result to obtain the projection of $v$ onto the set $\{z\mid f_{k}(z)\leq d\}$.
@@ 3.2 Optimality conditions
We characterize the optimality conditions of the projection problem; these conditions will be used in §4.3 to prove the correctness of the algorithm. Let $\mathcal{A}=\{a\in\{0,1\}^{m}\mid\mathbf{1}^{T}a=k\}$ be the set of all $\binom{m}{k}$ $k$-hot indicator vectors. Since $f_{k}(z)=\max_{a\in\mathcal{A}}a^{T}z$, the constraint $f_{k}(z)\leq d$ is equivalent to $a^{T}z\leq d$ for all $a\in\mathcal{A}$. We can therefore rewrite problem (11) in the equivalent form
[…]
where the subscript $i$ enumerates the elements of $\mathcal{A}$.
@@
The optimality conditions motivate an algorithm that solves the KKT system. We will define $z=v-\Delta$, where $\Delta$ is the vector we remove from $v$ to get the projection. The KKT conditions above imply that $\Delta$ is a nonnegative combination of the vectors $a_{i}$ for which $a_{i}^{T}z=d$, i.e., the indicators of the largest $k$ entries of $z$. (Note that there are potentially several such vectors in the presence of ties: consider the largest 2 entries in the vector $(2,1,1,0)$.) Thus, if we can construct a vector $\Delta$ that is a nonnegative combination of largest-$k$ indicator vectors, such that $f_{k}(v-\Delta)=d$, then we have found the projection $z=v-\Delta$.
@@ 4.1 Decrease step
is the set of $k$-hot indicator vectors whose first $n_{u}$ entries are equal to $1$ and all $1^{\prime}s$ occurring in the first $n_{u}+n_{t}$ entries.
@@ 4.2 Complexity
At each step, $n_{u}$ either decreases by $1$, or remains the same. At each step, $n_{t}$ increases by $1$. If the algorithm hasn’t terminated after $m$ steps, then there are $m$ tied entries, thus the subsequent decrease step will reduce the sum to the desired value, and the algorithm will terminate. Each step of the algorithm has constant complexity, and thus the algorithm (with sorted input) has complexity $O(m)$. Therefore, including the sorting and unsorting, the total complexity of the algorithm is $O(m\log m)$.
@@ 5 Numerical experiments
We compare our CVaR projection algorithm and our CVQP solver with two general-purpose solvers, Mosek and Clarabel. We first benchmark the projection algorithm on synthetic instances, then benchmark the CVQP solver on portfolio optimization and quantile regression problems. Code is available at https://github.com/cvxgrp/cvqp.
@@ 5.1 Experimental setup
All experiments were run on a Google Cloud n1-highmem-32 instance (32 vCPUs, 208 GB RAM), with a two-hour time limit per solve. For our CVQP solver we use ADMM penalty parameter $\rho^{(0)}=10^{-2}$, over-relaxation parameter $\alpha=1.7$, and update $\rho$ adaptively with parameters $\mu=10$ and $\tau=2$. We stop when primal and dual residuals satisfy absolute tolerance $10^{-4}$ and relative tolerance $10^{-3}$. Mosek and Clarabel are called through CVXPY with default settings. When we tighten our tolerances to $10^{-6}$, objective values agree with both solvers to about four significant figures.
@@ 5.2 CVaR projection
We generate random CVaR projection problems with vectors $v\in{\mbox{\bf R}}^{m}$ having entries uniformly distributed on $[0,1]$, with $k=(1-\beta)m$ and $\beta=0.95$. For each $v$, we set the constraint threshold to $d=\eta f_{k}(v)$ in problem (11), where $\eta\in(0,1)$ controls problem difficulty (larger $\eta$ gives easier problems), and project $v$ onto the convex set $\{z\mid f_{k}(z)\leq d\}$.